If the supply of corn is 𝑄n = b + 0.2p and the supply by the rest of the word is 𝑄r = e + 0.8p, what is the world supply.? What is the world inverse supply?



1
Expert's answer
2022-03-23T16:22:16-0400

If the supply of corn is Qn= b+ 0.2p

Supply by the rest of the world Qr=e+0.8p,

Then, the world supply is Qw = Qn + Qr = e + b +(0.2 + 0.8)*p

Therefore, world inverse supply is

P = Qw -(b+e)